# SPARTUN3D: Situated Spatial Understanding of 3D World in Large Language Models

### v1

@@ 1 Introduction
Recent advances in large language models (LLMs) have demonstrated their remarkable reasoning and communication capabilities across various tasks and modalities (Achiam et al., 2023; Alayrac et al., 2022; Zhang et al., 2023a; Rubenstein et al., 2023). Building on these breakthroughs, there has been a growing interest in extending LLMs to the 3D world (3D-based LLMs) (Hong et al., 2023; Huang et al., 2023; Chen et al., 2024; Zhen et al., 2024; Wang et al., 2023a). Existing studies mainly focus on integrating various 3D scene representations into LLMs, enabling the models to perform 3D grounding and spatial reasoning through natural language. For example, 3D-LLM (Hong et al., 2023) utilizes multi-view images to represent 3D scenes, pioneering a new direction in this field, while LEO (Huang et al., 2023) further pushes the boundary by directly injecting 3D point clouds into LLMs, aiming to develop a generalist embodied agent capable of 3D grounding, embodied reasoning, and action planning.
Despite the promising progress, current 3D-based LLMs still fall short in situated understanding, a fundamental capability for completing embodied tasks, such as Embodied Question Answering (Das et al., 2018), Vision and Language Navigation (Anderson et al., 2018; Wang et al., 2023b; Zhang et al., 2024c), robotic manipulation (Shridhar et al., 2022), and many others. Situated understanding refers to the ability to interpret and reason about a 3D scene from a dynamic egocentric perspective, where the agent must continuously adjust understanding based on its changing position and evolving environment around it. This capability is crucial because an agent’s reasoning and response to the same question can vary depending on its current situation. For example, as shown in Fig 1, given the same question “What should you do to wash hands?”, the agent might need to answer “use the sink on the right” or “use the sink on the left”, depending on the agent’s current perspective and location relative to the “sink”.
@@
Figure 2: Examples of two tasks in Spartun3D Benchmark. The green bounding box and arrow in the 3D scene demonstrate the standing point and orientation.
We conduct extensive experiments across a variety of tasks, including Spartun3D, Situated QA (Ma et al., 2022), and MP3D Nav (Savva et al., 2019). Our results demonstrate that our model, trained on Spartun3D, exhibits strong generalization to other tasks, specifically in zero-shot settings, highlighting the effectiveness of our proposed dataset. Additionally, Spartun3D-LLM outperforms the baseline on nearly all tasks, indicating that incorporating direct text supervision improves the model’s spatial understanding ability. This is further supported by our observation that the explicit alignment module improves the generation of more fine-grained, context-aware spatial information.
@@ 2 Related Work
Situated Scene Understanding
is essential for various embodied tasks, including embodied QA (Das et al., 2018; Wijmans et al., 2019; Majumdar et al., 2024), vision-and-language navigation (Anderson et al., 2018; Savva et al., 2019; Gu et al., 2022; Zhang & Kordjamshidi, 2022b; Zhang et al., 2021; Hong et al., 2022; Zhang & Kordjamshidi, 2022a; 2024) and robotic manipulation (Shridhar et al., 2022; Jiang et al., 2022; Driess et al., 2023). Recently, SQA3D (Ma et al., 2022) introduces a human-annotated dataset where the model generates answers based on questions and given situations. SIG3D (Man et al., 2024) highlights the situated awareness via a situation-grounded 3D VL reasoning architecture. While SQA3D is crucial in 3D vision-language learning (Zhu et al., 2023; Huang et al., 2023; Hong et al., 2023), its reliance on human annotations makes it costly and difficult to scale for the large-scale training required by 3D-based LLMs. In contrast, we use the LLM to design an automated pipeline to generate situated questions and answers, enhancing both the scalability and diversity in data collection procedure.
Grounded 3D Scene Understanding.
Compared to 2D vision-language tasks (Lin et al., 2014; Antol et al., 2015; Zhang et al., 2024a), 3D scenes introduce additional dimensions of knowledge that is more challenging to align with text modalities. Early studies in this area primarily focused on grounding language to individual objects within 3D environments (Achlioptas et al., 2019; 2020; Chen et al., 2020; 2019; 2021). Recently, 3D Vista (Zhu et al., 2023) proposes a pre-trained VL Transformer for 3D vision and text alignment, and a few works utilize LLMs in understanding 3D scenes (Hong et al., 2023; Huang et al., 2023; Chen et al., 2024; Zhen et al., 2024; Yang et al., 2024). However, these works are less effective in situated understanding tasks that require generating answers from the agent’s dynamic perspectives, which is the key challenge that this work is tackling.
@@ 3.1 Situation Design
Standing Point and Orientation.
We begin with determining the agent’s standing point and orientation within the 3D scene. Our approach is to place the agent beside an object, ensuring a clear reference for orientation when interacting with the environment. Specifically, we project all objects from 3D space onto a 2D plane, focusing only on the x and y coordinates to construct a bird-eye-view of the scene. From this projected 2D scene, we randomly select an object from the set of segmented objects within the 3D scene. To ensure the constructed situation remains realistic, we exclude objects that are positioned too high to avoid unnatural situations like “standing beside the lamp on the ceiling”. As a result, we limit the selection to objects whose z-axis is below the average height of all objects in the scene. Then, we choose a midpoint from two sides of the selected object’s bounding box that are closest to the center of the scene, as shown in Fig. 4. By prioritizing the side closest to the center, we minimize this risk and keep the agent within the scene’s boundaries. Finally, the selected midpoint will be used as the agent’s standing point. In addition, we need to determine the agent’s orientation. We assume the agent’s orientation is always facing forward to the center of the selected object. This guarantees that the selected object remains within the agent’s field of view.
Pivot and Referent Object.
Once the agent’s standing point and orientation are determined, we refer to the object that the agent stands beside as the “pivot object”, and other objects surrounding the pivot object are potential referent objects. A referent object is then randomly selected, and its relative position (left/right/front/backward), with respect to the agent’s standing point and orientation, is used to generate the description of the situation.
@@ 3.3 LLMs prompting
Situated Captioning
is our newly introduced task, aiming to generate brief situated descriptions of the surrounding objects as the agent performs a $360^{\circ}$ clockwise rotation starting from its standing point and orientation. The motivation for introducing this task stems from its crucial role in embodied tasks, such as navigation, where the agent must interpret and reason about its environment from $360^{\circ}$ panoramic views to make decisions about movement and interaction (Zhang & Kordjamshidi, 2023; Zhou et al., 2024; Zhang et al., 2024b). Therefore, we guide GPT-4o to generate descriptions progressively, starting from lower rotation angles and moving toward higher angles in each direction. However, in practice, certain directions may not contain any objects, leading GPT-4o to omit those areas. To address this, we employ a template-based approach that enforces the model to generate descriptions for every direction. If there is no object in a certain direction, the description should mention the absence of objects. Each direction description is then combined with a sequence of directions: front, right, backward, and left, following a clockwise rotation.
Situated QA.
We design three types of questions for the Situated QA task, each targeting a different aspect of spatial reasoning for embodied agents. Unlike previous works that rely on a single generic prompt for all question types, we develop tailored prompting strategies for each question type, encouraging LLM to generate QA pairs focusing on different levels of reasoning.
@@ 3.4 Dataset Statistics and Quality Evaluation
Tasks # of Examples Train/Test Cord-prompt Spa-prompt
Captioning $\sim 10$K $8,367$/$1,350$ $27\%$ $86\%$
Attr. & Rel. $\sim 62$K $61,254$/$8,168$ $57\%$ $90\%$
Affordance $\sim 40$K $35,070$/$5,017$ $40\%$ $87\%$
Planning $\sim 21$K $19,434$/$2,819$ $33\%$ $87\%$
Figure 6: Spartun3D-LLM Model Architecture.
To evaluate the quality of Spartun3D, we randomly sampled approximately $50$ examples for each task and manually assessed their validity. Each example is evaluated based on whether the generated captions or answers are accurate given the 3D scene and the agent’s situation. We experiment with two types of prompts for representing spatial information to prompt GPT-4O : Cord-prompt, which consists of object center coordinates, standing point, orientation, and instructions for calculating distances and rotation angles, and Spa-prompt, consisting of the calculated angles and distance based on the approaches we described in Sec. 3.3. An example of each type of prompt can be found in Tab. 10 in the Appendix. Tab. 1 shows the percentage for valid outputs based on different prompts. The results indicate that Cord-prompt yields unsatisfactory results, revealing that LLMs lack strong 3D reasoning capabilities when tasked with interpreting raw spatial coordinates. Spa-prompt helps improve the quality of the generated dataset by providing more straightforward spatial information. The high valid percentage based on Spa-prompt also demonstrates the high quality of our dataset.
@@ 4.1 Background
Backbone.
LEO (Huang et al., 2023) takes the text, 2D image (optional), and 3D point clouds as input and formulate comprehensive 3D tasks as autoregressive sequence generation. Specifically, data from different modalities are converted into a sequence of tokens as input to the LLM. The text tokens include system messages (e.g., “You are an AI visual assistant situated in a 3D scene.”), situations, and questions. These tokens are then embedded into vector representations using an embedding look-up table. For 3D point clouds, LEO first applies segmentation masks to extract the point clouds of individual objects in the 3D scenes. Then, the sampled points of each object are input into a object-centric point cloud encoder, PointNet++ (Qi et al., 2017) pre-trained on ScanNet (Dai et al., 2017), to obtain the object-level representations.
@@ 4.2 Situated Spatial Alignment Module
Situated Textual Descriptions.
For each object, we construct a comprehensive situated textual description based on a template that captures the object’s name, attributes, and spatial relations with nearby objects, as “Stand besides {object name} and facing the center of the {object name}, in front, there are {a list of nearby objects}; on the right, …; behind …; and on the left…”. The object’s attributes are also considered (e.g., “white chair”). We consider up to five objects per direction. If no object is present in a specific direction, the description explicitly states this,yueensuring to provide complete information about the 3D environment.
3D Object-Text Alignment.
Inspired by the success of 2D Visual-Language models, which effectively leverage semantically aligned text and visual features to excel in downstream tasks (Radford et al., 2021; Li et al., 2022; 2023; Wang et al., 2022), we aim to enhance the 3D visual representations so that they can better encode the situated spatial information and effectively align with the textual descriptions. Specifically, we introduce a 3D object-text alignment loss to guide the learning process of point cloud encoders within 3D-based LLMs, leveraging the robust language representations captured by pre-trained text encoders. We experiment with various text encoders, and CLIP achieved the best performance. For more details, please refer to the Sec. A.3 in the Appendix.
@@ 5.1 Experimental Setup
To demonstrate the effectiveness of our proposed Spartun3D-LLM, we conduct experiments on two situated understanding datasets, including Spartun3D and SQA3D (Ma et al., 2022). For SQA3D, we evaluate under two conditions: object proposals in 3D are derived either from Mask3D (Schult et al., 2023) or ground-truth annotations. Also, we assess the transferability of our method on the navigation task using MP3D ObjNav (Savva et al., 2019). Following LEO (Huang et al., 2023), we report the performance using standard generation metrics, including CIDEr, METEOR, BLEU-4, and ROUGE_L, sentence similarity (Reimers, 2019) for captioning task. For SQA3D and situated QA tasks of questions about attributes and relations, we also report an additional metric of exact-match accuracy. More details about metrics are introduced in Appendix A.1. We also provide implementation details in Sec. A.2. We leverage LEO as baseline. Since the training stage in LEO has covered most of the evaluation tasks, we experiment with models initialized from scratch to ensure a fair comparison in the zero-shot setting. For other settings, we report the performance of models initialized both from scratch and from the instruction-tuned LEO. To distinguish between the two, models initialized from the instruction-tuned LEO are marked with an asterisk ().
Methods Mask3D (Schult et al., 2023) GT
C M R EM C M R EM
zero-shot LEO (Huang et al., 2023) $14.2$ $6.4$ $8.2$ $12.4$ $15.3$ $6.7$ $8.6$ $13.9$
LEO+Spartun3D $82.3$ $14.2$ $32.8$ $34.7$ $83.1$ $15.2$ $33.7$ $35.9$
Spartun3D-LLM $83.5$ $15.7$ $34.7$ $36.2$ $85.6$ $16.6$ $35.8$ $37.1$
Fine-tune 3D-Vista (Zhu et al., 2023) - - - $48.5$
3D-LLM (Hong et al., 2023) - - - $50.2$
LEO (Huang et al., 2023) $132.0$ $33.0$ $49.2$ $52.4$ $132.3$ $34.3$ $51.4$ $52.5$
LEO*+Spartun3D $134.0$ $34.6$ $52.2$ $53.5$ $135.3$ $34.2$ $52.1$ $54.2$
Spartun3D-LLM* $\mathbf{138.2}$ $\mathbf{35.3}$ $\mathbf{53.4}$ $\mathbf{54.8}$ $\mathbf{138.3}$ $\mathbf{35.4}$ $\mathbf{53.7}$ $\mathbf{55.0}$

### v2

@@ 1 Introduction
Recent advances in large language models (LLMs) have demonstrated their remarkable reasoning and communication capabilities across various tasks and modalities (Achiam et al., 2023; Alayrac et al., 2022; Zhang et al., 2023a; Rubenstein et al., 2023). Building on these breakthroughs, there has been a growing interest in extending LLMs to the 3D world (3D-based LLMs) (Hong et al., 2023; Huang et al., 2023; Chen et al., 2024; Zhen et al., 2024; Wang et al., 2023). Existing studies mainly focus on integrating various 3D scene representations into LLMs, enabling the models to perform 3D grounding and spatial reasoning through natural language. For example, 3D-LLM (Hong et al., 2023) utilizes multi-view images to represent 3D scenes, pioneering a new direction in this field, while LEO (Huang et al., 2023) further pushes the boundary by directly injecting 3D point clouds into LLMs, aiming to develop a generalist embodied agent capable of 3D grounding, embodied reasoning, and action planning.
Despite the promising progress, current 3D-based LLMs still fall short in situated understanding, a fundamental capability for completing embodied tasks, such as Embodied Question Answering (Das et al., 2018), Vision and Language Navigation (Anderson et al., 2018; Zhang et al., 2024c), robotic manipulation (Shridhar et al., 2022), and many others. Situated understanding refers to the ability to interpret and reason about a 3D scene from a dynamic egocentric perspective, where the agent must continuously adjust understanding based on its changing position and evolving environment around it. This capability is crucial because an agent’s reasoning and response to the same question can vary depending on its current situation. For example, as shown in Fig 1, given the same question “What should you do to wash hands?”, the agent might need to answer “use the sink on the left/right” based on the agent’s current perspective and location relative to the “sink”.
@@
Figure 2: Examples of Spartun3D. Green box and arrow show the standing point and orientation.
We conduct extensive experiments across a variety of tasks, including Spartun3D, SQA3D, and MP3D Nav (Savva et al., 2019). Our results demonstrate that our model, trained on Spartun3D, exhibits strong generalization to other tasks, specifically in zero-shot settings, highlighting the effectiveness of our proposed dataset. Additionally, Spartun3D-LLM outperforms the baseline on nearly all tasks, indicating that incorporating direct text supervision improves the model’s spatial understanding ability. This is further supported by our observation that the explicit alignment module improves the generation of more fine-grained, context-aware spatial information.
@@ 2 Related Work
Situated Scene Understanding is essential for various embodied tasks, including embodied QA (Das et al., 2018; Wijmans et al., 2019), vision-and-language navigation (Anderson et al., 2018; Zhang et al., 2021; Zhang & Kordjamshidi, 2022a; b; 2024) and robotic manipulation (Shridhar et al., 2022; Jiang et al., 2022; Driess et al., 2023). Recently, SQA3D (Ma et al., 2022) introduces a human-annotated dataset where the model generates answers based on questions and given situations. SIG3D (Man et al., 2024) highlights the situated awareness via a situation-grounded 3D VL reasoning architecture. While SQA3D is crucial in 3D vision-language learning (Zhu et al., 2023; Huang et al., 2023), its reliance on human annotations makes it costly and difficult to scale for the large-scale training required by 3D-based LLMs. In contrast, we use the LLM to design an automated pipeline to generate situated questions and answers, enhancing both the scalability and diversity in data collection procedure.
Grounded 3D Scene Understanding. Compared to 2D vision-language tasks (Song et al., 2024; Antol et al., 2015; Zhang et al., 2024a; Guo et al., 2024a; b; Xu et al., 2024; Wang et al., 2024; Qi et al., 2024; Guo et al., 2025), 3D scenes introduce additional dimensions of knowledge that is more challenging to align with text modalities. Early studies in this area primarily focused on grounding language to individual objects within 3D environments (Achlioptas et al., 2019; 2020; Chen et al., 2020; 2019; 2021). Recently, 3D Vista (Zhu et al., 2023) proposes a pre-trained VL Transformer for 3D vision and text alignment, and a few works utilize LLMs in understanding 3D scenes (Hong et al., 2023; Huang et al., 2023; Chen et al., 2024; Zhen et al., 2024; Yang et al., 2024). However, these works are less effective in situated understanding tasks that require generating answers from the agent’s dynamic perspectives. MSQA (Linghu et al., 2025) is a concurrent work that also explore situated understanding of 3D-based LLMs. However, our approach differs in both dataset construction and model design. While MSQA focuses on leveraging images to assist in describing a situation, our method emphasizes directly understanding the situation from textual descriptions, leading to distinct contributions in the field.
@@ 3.1 Situation Design
Standing Point and Orientation. We begin with determining the agent’s standing point and orientation within the 3D scene. Our approach is to place the agent beside an object, ensuring a clear reference for orientation when interacting with the environment. Specifically, we project all objects from 3D space onto a 2D plane, focusing only on the x and y coordinates to construct a bird-eye-view of the scene. From this projected 2D scene, we randomly select an object from the set of segmented objects within the 3D scene. To ensure the constructed situation remains realistic, we exclude objects that are positioned too high to avoid unnatural situations like “standing beside the lamp on the ceiling”. As a result, we limit the selection to objects whose z-axis is below the average height of all objects in the scene. Then, we choose a midpoint from two sides of the selected object’s bounding box that are closest to the center of the scene, as shown in Fig. 4. By prioritizing the side closest to the center, we minimize this risk and keep the agent within the scene’s boundaries. Finally, the selected midpoint will be used as the agent’s standing point. In addition, we need to determine the agent’s orientation. We assume the agent’s orientation is always facing forward to the center of the selected object. This guarantees that the selected object remains within the agent’s field of view.
Pivot and Referent Object. Once the agent’s standing point and orientation are determined, we refer to the object that the agent stands beside as the “pivot object”, and other objects surrounding the pivot object are potential referent objects. A referent object is then randomly selected, and its relative position (left/right/front/backward), with respect to the agent’s standing point and orientation, is used to generate the description of the situation.
@@ 3.3 LLMs prompting
Situated Captioning is our newly introduced task, aiming to generate brief situated descriptions of the surrounding objects as the agent performs a $360^{\circ}$ clockwise rotation starting from its standing point and orientation. The motivation for introducing this task stems from its crucial role in embodied tasks, such as navigation, where the agent must interpret and reason about its environment from $360^{\circ}$ panoramic views to make decisions about movement and interaction (Zhou et al., 2024; Zhang & Kordjamshidi, 2023; Zhang et al., 2024b). Therefore, we guide GPT-4o to generate descriptions progressively, starting from lower rotation angles and moving toward higher angles in each direction.
Situated QA. We design three types of questions for the Situated QA task, each targeting a different aspect of spatial reasoning for embodied agents. Unlike previous works that rely on a single generic prompt for all question types, we develop tailored prompting strategies for each question type, encouraging LLM to generate QA pairs focusing on different levels of reasoning.
@@ 3.4 Dataset Statistics and Quality Evaluation
Tasks # of Examples Train/Test
Captioning $\sim 10$K $8,367$/$1,350$
Attr. & Rel. $\sim 62$K $61,254$/$8,168$
Affordance $\sim 40$K $35,070$/$5,017$
Planning $\sim 21$K $19,434$/$2,819$
Figure 6: Human Evaluation of Spartun3D.
We conduct human evaluation to assess the quality of Spartun3D, introducing scores based on two key criteria: language naturalness, which evaluates whether the text reads as if it were naturally written by a human, and spatial fidelity, which ensures that the data accurately reflects the 3D scene with correct spatial relationships. Detailed explanations of the criteria are in Sec. A.1. Each criterion is rated on a scale from $1$ to $5$, and the average of these two scores is the overall human score. We randomly select $50$ examples from each task and compute human scores of situation, question, and answer, respectively. As shown in Fig. 6 (a), the average scores align with the complexity of each task, with relatively lower scores for captioning and planning tasks. To evaluate how our generated data compares to human-annotated data, we sampled $50$ examples from SQA3D and combined them with our dataset. Our data shows a similar trend in human evaluation results across different question types as observed in SQA3D (shown in Fig. 6 (b)). We finally evaluate how different prompting strategies influence the quality of the data. We experiment with two types of prompts for representing spatial information to prompt GPT-4o: Cord-prompt, which consists of object center coordinates, standing point, orientation, and instructions for calculating distances and rotation angles, and Spa-prompt, consisting of the calculated angles and distance based on the approaches we described in Sec. 3.3. An example of each type of prompt can be found in Tab. 11. Fig.6 (c) shows the percentage of examples with high human scores ($\geq 4$) for each prompt across tasks. The results indicate that Cord-prompt yields unsatisfactory results, revealing that LLMs lack strong 3D spatial reasoning when interpreting raw spatial coordinates, which is consistent with their struggles in spatial reasoning across text and 2D images (Zhang et al., 2024d; Premsri & Kordjamshidi, 2024; 2025). Our Spa-prompt significantly improves the quality of the generated dataset by providing qualitative spatial relations (e.g. distance, direction).
@@ 4.1 Background
Figure 7: Spartun3D-LLM Model Architecture.
Backbone. LEO takes text, 2D image (optional), and 3D point clouds as input and formulate comprehensive 3D tasks as autoregressive sequence generation. Specifically, data from different modalities are converted into a sequence of tokens as input to the LLM. The text tokens include system messages (e.g., “You are an AI visual assistant situated in a 3D scene.”), situations, and questions. These tokens are then embedded into vector representations using an embedding look-up table. For 3D point clouds, LEO first applies segmentation masks to extract the point clouds of individual objects in the 3D scenes. Then, the sampled points of each object are input into a object-centric point cloud encoder, PointNet++ (Qi et al., 2017) pre-trained on ScanNet (Dai et al., 2017), to obtain the object-level representations.
@@ 4.2 Situated Spatial Alignment Module
Situated Textual Descriptions. For each object, we construct a comprehensive situated textual description based on a template that captures the object’s name, attributes, and spatial relations with nearby objects, as “Stand besides {object name} and facing the center of the {object name}, in front, there are {a list of nearby objects}; on the right, …; behind …; and on the left…”. The object’s attributes are also considered (e.g., “white chair”). We consider up to five objects per direction. If no object is present in a specific direction, the description explicitly states this, ensuring to provide complete information about the 3D environment.
3D Object-Text Alignment. Inspired by the success of 2D Visual-Language models, which effectively leverage semantically aligned text and visual features to excel in downstream tasks (Radford et al., 2021; Li et al., 2022; 2023), we aim to enhance the 3D visual representations so that they can better encode the situated spatial information and effectively align with the textual descriptions. Specifically, we introduce a 3D object-text alignment loss to guide the learning process of point cloud encoders within 3D-based LLMs, leveraging the robust language representations captured by pre-trained text encoders. We experiment with various text encoders, and CLIP achieved the best performance. For more details, please refer to the Sec. A.4 in the Appendix.
@@ C B-4 M R EM C B-4 M R S C B-4 M R S
# Methods Mask3D (Schult et al., 2023) GT
C M R EM C M R EM
Zero-shot $1$ LEO (Huang et al., 2023) $14.2$ $6.4$ $8.2$ $12.4$ $15.3$ $6.7$ $8.6$ $13.9$
$2$ LEO+Spartun3D $82.3$ $14.2$ $32.8$ $34.7$ $83.1$ $15.2$ $33.7$ $35.9$
$3$ Spartun3D-LLM $83.5$ $15.7$ $34.7$ $36.2$ $85.6$ $16.6$ $35.8$ $37.1$
Fine-tune $4$ 3D-Vista (Zhu et al., 2023) - - - $48.5$
$5$ 3D-LLM (Hong et al., 2023) - - - $50.2$
$6$ LEO (Huang et al., 2023) $132.0$ $33.0$ $49.2$ $52.4$ $132.3$ $34.3$ $51.4$ $52.5$
$7$ LEO*+Spartun3D $134.0$ $34.6$ $52.2$ $53.5$ $135.3$ $34.2$ $52.1$ $54.2$
$8$ Spartun3D-LLM* $\mathbf{138.2}$ $\mathbf{35.3}$ $\mathbf{53.4}$ $\mathbf{54.8}$ $\mathbf{138.3}$ $\mathbf{35.4}$ $\mathbf{53.7}$ $\mathbf{55.0}$
@@ 5.1 Experimental Setup
To demonstrate the effectiveness of our proposed Spartun3D-LLM, we conduct experiments on two situated understanding datasets: Spartun3D [^1] https://github.com/zhangyuejoslin/Spartun3D and SQA3D. For SQA3D, we evaluate under two conditions: object proposals in 3D are derived either from Mask3D (Schult et al., 2023) or ground-truth annotations. Also, we assess the transferability of our method on the navigation task using MP3D ObjNav (Savva et al., 2019). Following LEO (Huang et al., 2023), we report the performance using standard generation metrics, including CIDEr, METEOR, BLEU-4, and ROUGE_L, sentence similarity (Reimers, 2019) for captioning task. For SQA3D and situated QA tasks of questions about attributes and relations, we also report an additional metric of exact-match accuracy. More details about metrics are introduced in Appendix A.2.1. We also provide implementation details in Sec. A.3. We leverage LEO as baseline. Since the training stage in LEO has covered most of the evaluation tasks, we experiment with models initialized from scratch to ensure a fair comparison in the zero-shot setting. For other settings, we report the performance of models initialized both from scratch and from the instruction-tuned LEO. To distinguish between the two, models initialized from the instruction-tuned LEO are marked with an asterisk ().